OpenStudy (anonymous):

what are the solutions of the quadratic equation? 5x^2+37x+14=0

OpenStudy (anonymous):

(-b plus or minus sqrt(b^2-4ac) ) / 2a

OpenStudy (anonymous):

I tried doing that but I get decimals

OpenStudy (anonymous):

-7,-2/5

OpenStudy (anonymous):

\[\left\{ -b \left(\begin{matrix}+ \\ -\end{matrix}\right)\sqrt{b ^{2}-4ac} \right\} \div \left\{ 2a \right\}\]

OpenStudy (anonymous):

@devinGT96.. stop spamming everyone.

OpenStudy (anonymous):

I need the steps to do so (I have a math test tomorrow and need to how to do this) DevinGT96

OpenStudy (anonymous):

yeah don't listen to her. are you sure your doing it right? label your variables a b and c and plug in the numbers to the left of the variables into the quadratic equation

OpenStudy (anonymous):

Yeah I am super lost . . .

OpenStudy (anonymous):

here ...